# MGE-LDM: Joint Latent Diffusion for Simultaneous Music Generation and Source Extraction

### v1

@@ Abstract
We present MGE-LDM, a unified latent diffusion framework for simultaneous music generation, source imputation, and query-driven source separation. Unlike prior approaches constrained to fixed instrument classes, MGE-LDM learns a joint distribution over full mixtures, submixtures, and individual stems within a single compact latent diffusion model. At inference, MGE-LDM enables (1) complete mixture generation, (2) partial generation (i.e., source imputation), and (3) text-conditioned extraction of arbitrary sources. By formulating both separation and imputation as conditional inpainting tasks in the latent space, our approach supports flexible, class-agnostic manipulation of arbitrary instrument sources. Notably, MGE-LDM can be trained jointly across heterogeneous multi-track datasets (e.g., Slakh2100, MUSDB18, MoisesDB) without relying on predefined instrument categories. Audio samples are available at our project page  [^1] https://yoongi43.github.io/MGELDM_Samples/.
@@ 1 Introduction
Recent advances in generative modeling have significantly accelerated progress in music audio synthesis, inspired by breakthroughs in the language and vision domains. Early autoregressive models such as WaveNet [1] demonstrated the feasibility of end-to-end waveform generation. Since then, two dominant approaches have emerged: (1) discrete-token models, which compress raw audio into quantized codes [2, 3, 4, 5] for sequence modeling [6, 7, 8]; and (2) diffusion-based models [9, 10, 11], which synthesize audio by reversing a noise corruption process in the waveform domain [12, 13]. Building on this foundation, latent diffusion models (LDMs) [14]—which operate in a compressed latent space—have delivered substantial gains in synthesis quality and have been successfully applied to music generation tasks [15, 16, 17, 18].
Despite this progress, most music generation models produce a single, mixed waveform, lacking access to the individual instrument stems required for remixing, adaptive arrangement, or downstream production tasks. To recover these components, audio source separation techniques aim to decompose a mixture into its constituent tracks. Discriminative approaches [19, 20, 21, 22] learn to directly regress each source from the input mixture, achieving strong performance. In contrast, generative separation models sample individual sources from a learned prior [23, 24, 25, 26]. Recently, diffusion-based methods have demonstrated strong performance not only in speech separation [27, 28], but also in speech enhancement tasks [29, 30, 31, 32], highlighting their potential for flexible, high-quality source recovery across audio domains.
@@
To address these limitations, we introduce MGE-LDM, a class-agnostic latent diffusion framework that jointly unifies music generation, partial generation, and arbitrary source extraction. Our approach models three interrelated latent variables—mixture, submixture, and source—within a single diffusion backbone.
Given a waveform mixture $x^{(m)}$ and a chosen source $x^{(s)}$, we define the submixture as $x^{(u)}=x^{(m)}-x^{(s)}$. A pretrained encoder $E$ maps each waveform to its latent representations:
@@
At inference time, diffusion-based inpainting – conditioning on a subset of $\{z^{(m)},z^{(u)},z^{(s)}\}$ – enables class-agnostic partial generation and text-driven source extraction, without relying on fixed instrument vocabularies or linear mixing assumptions. A detailed formulation appears in Section 3.
@@
In contrast, our three-track formulation—mixture, submixture, and individual source—naturally accommodates such aggregated labels, allowing them to be interpreted as submixtures and thereby incorporated during joint training alongside fully separated stems. This dataset flexibility simplifies data collection and enhances model robustness across heterogeneous source structures.
@@ 2 Related Work
Audio Source Separation and Extraction. Audio source separation aims to decompose a polyphonic mixture into its constituent tracks, while source extraction focuses on isolating a particular target sound, often guided by metadata, text prompts, or reference examples. Two dominant paradigms have emerged: discriminative models learn a direct mapping from the input mixture to each target stem via regression losses, either in the waveform domain or on spectrogram representations [40, 19, 41, 20, 21, 22]. In contrast, generative approaches learn probabilistic priors over source distributions and recover individual stems via sampling [23, 24, 26, 25, 42].
@@
Parallel efforts introduced discrete-token models, where audio is encoded into compact code sequences using vector quantization (e.g., VQ-VAE [2]). Jukebox [6] models long-range dependencies over codes using Transformers [59], while recent systems enhance fidelity through residual quantization [3, 4, 5] and hierarchical token modeling, where coarse-to-fine code representations are generated over multiple levels [60, 61, 7, 62]. MusicGen [8] improves decoding efficiency with delayed-token generation, and Instruct-MusicGen [63] extends it for targeted editing via instruction-tuned prompts. Concurrently, token-based masked generative modeling techniques—originally developed for the vision domain [64]—have been extended to audio, enabling efficient non-autoregressive synthesis and precise spectrogram inpainting [65, 66].
@@
Multi-Track Music Audio Modeling. Recent studies model multi-track music as a structured composition of interdependent stems. StemGen [74] employs an iterative, non-autoregressive transformer over discrete tokens to generate stems conditioned on text prompts. Jen-1 Composer [75] applies latent diffusion to jointly model four canonical stems (bass, drums, instrument, melody), producing coherent multi-track compositions. MusicGen-Stem [74] combines per-stem vector quantization with an autoregressive decoder to synthesize bass, drums, and aggregated other components, and supports mixture-conditioned accompaniment generation.
Other work explores joint modeling of synthesis and decomposition within a single diffusion backbone. Multi-Source Diffusion Models (MSDM) [33] model a fixed set of stems (bass, drums, guitar, and piano) within a shared diffusion framework, relying on an additive mixture assumption and a Dirac delta-based posterior sampler, following the EDM formulation for ODE-based sampling [76]. This line of work has since been extended in GMSDI [34], MSG-LD [35], and others [39, 77]. GMSDI enables variable-stem modeling and text-based conditioning but remains grounded in waveform-space additive mixing. MSG-LD adapts latent diffusion for four-stem modeling and classifier-free guidance [78], though it still assumes fixed instrument classes.
@@ 3 Method
Figure 1 provides an overview of the proposed MGE-LDM pipeline. We begin by defining the construction of training triplets and then describe how they are used for joint diffusion-based training and inpainting-based inference.
@@ 3.2 Latent Diffusion Training with Three-Track Embeddings
where $z^{(k)}_{0}\in\mathbb{R}^{C\times L}$ are (clean) track embeddings, with $k\in K=\{m,u,s\}$ denoting the track type – mixture, submixture, and source, respectively.
We aim to estimate the score $\nabla_{\textbf{z}_{\tau}}\log q_{\tau}(\textbf{z}_{\tau})$ of the perturbed latent $\textbf{z}_{\tau}$ across continuous noise levels $\tau\in\left[\tau_{\text{min}},1\right]$. To do so, we perturb $\textbf{z}_{0}$ with Gaussian noise according to:
@@
where the noise scaling coefficients are defined as:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{\textbf{z}_{0},\boldsymbol{\epsilon},\tau}\left||f_{\theta}(\textbf{z}_{\tau},\tau,\mathbf{c})-\boldsymbol{v}_{\tau}\right||_{2}^{2},\quad\boldsymbol{v}_{\tau}=\frac{\partial\textbf{z}_{\tau}}{\phi_{\tau}}=\alpha_{\tau}\boldsymbol{\epsilon}-\beta_{\tau}\textbf{z}_{0}.$ (4)
@@ 3.3 Inference via Conditional Sampling in Latent Space
Total Generation. Let $p_{\theta}(z^{(m)},z^{(u)},z^{(s)})$ denote the implicit model distribution whose score function is induced by the denoising network $f_{\theta}$. To synthesize a complete mixture, we condition only on the mixture prompt embedding $c^{(m)}$, or omit all conditions for unconditional generation. We sample mixture latent $\hat{z}^{(m)}$ as:
@@
Let ¯=$\bar{\mathcal{I}}=\{c_{1},...,c_{J}\}$ be an (ordered) set of CLAP-derived text embeddings, each corresponding to a target source description to be imputed. Let $x^{(u)}_{0}$ denote the waveform mixture of the observed sources and let $z^{(u)}_{0}=E(x^{(u)}_{0})$ be its latent representation. We initialize the submixture latent with $z^{(u)}_{0}$, and generate each missing source sequentially.
@@
After $J$ iterations, we obtain the full set of imputed sources $\{\hat{z}^{(s)}_{j}\}_{j=1}^{J}$, and reconstruct the final mixture waveform as:
@@ 3.4 Track-aware Inpainting Model with Adaptive Timesteps
where $\alpha_{\tau}$ and $\beta_{\tau}$ is cosine schedule as defined in Eq. (1), following [81]. For notational simplicity, we omit conditioning vectors $\mathbf{c}$ in this expression. A detailed derivation is provided in Appendix A.1.
Recently, region-aware adaptations of diffusion inpainting – such as spatially varying noise schedules [85] and per-pixel timestep conditioning in TD-Paint [86] – have demonstrated substantial improvements in semantic consistency by preserving fidelity in observed regions. Inspired by TD-Paint, we extend this idea to three-track music audio by assigning distinct timestep conditions to each track, thereby improving inpainting quality in the latent space.
We describe our track-wise adaptive timestep conditioned model using general notation. Let $K$ be a set of tracks, and let $N=|K|$ be the number of tracks. Define the clean latent tensor and corresponding noise levels as:
@@
and extend this notation to the cosine noise schedule terms as:
@@
The denoiser $f_{\theta}(\textbf{z}_{\boldsymbol{\tau}},\boldsymbol{\tau})$ is trained to regress the velocity target under v-objective:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{\textbf{z}_{0},\boldsymbol{\epsilon},\boldsymbol{\tau}}||f_{\theta}(\textbf{z}_{\boldsymbol{\tau}},\boldsymbol{\tau})-\boldsymbol{v}_{\boldsymbol{\tau}}||_{2}^{2}.$ (12)
@@
During training, we first sample a noise level $\boldsymbol{\tau}\sim\mathcal{U}([\tau_{\text{min}},1])$, and set the per-track timestep vector $\boldsymbol{\tau}\in\mathbb{R}^{3}$ according to one of the following four patterns:
@@
[…]
For example when $(\tau_{m},\tau_{u},\tau_{s})=(0,\tau,\tau)$, the model is trained to approximate the gradient:
@@
which the joint-score formulation in Eq. (9) cannot compute in closed form. At inference time, we clamp observed tracks by setting their noise levels to zero, and apply standard reverse diffusion updates to the remaining (missing) tracks. Pseudocode and a detailed theoretical comparison with conventional inpainting methods are provided in Appendix B.
@@ 4 Experimental Setup
Baselines. We use two recent multi-track diffusion models – MSDM [33] and MSG-LD [35] – as baselines, both of which operate on a fixed set of stems: bass, drums, guitar, and piano. In addition to generative and inpainting performance, we assess source extraction capabilities against Demucs [41], which separates the mixture into bass, drums, other, and vocals stems, and AudioSep [52], which performs text-conditioned separation based on natural language queries.
@@
Implementation Details. Our models use the Stable Audio backbone [71], comprising an autoencoder and a DiT-based diffusion model. To better accommodate per-track variability in the joint latent space, we replace LayerNorm [87] with GroupNorm [88], using three groups to reflect the number of tracks.
@@
All of our models – except the one trained on the full dataset combination ($\textbf{S}_{A}\,$, $\textbf{S}_{B}\,$, $\textbf{M}_{u}\,$, $\textbf{M}_{o}\,$) – are trained for 200K iterations with a batch size of 64, using 16 kHz audio segments of 10.24 seconds. The full combination model is trained for 320K iterations with a batch size of 128. During sampling and inpainting, we apply classifier-free guidance (CFG) with a guidance scale of 2.0 and a per-track dropout probability of $p=0.1$. All diffusion-based samples – including those from baseline models – are generated using 250 inference steps. We adopt DDIM sampling [90] for all our models, while each baseline uses its originally proposed sampling method.
To support reproducibility, we plan to release all code and pretrained models upon acceptance.
5 Result
Table 1: Total generation results. Reported scores are FAD $\downarrow$, computed against mixture references from each test set. Values in parentheses indicate generation results conditioned on the prompt "The sound of the bass, drums, guitar, and piano", as detailed in Section 5.1.
@@ 5.1 Music Generation
To assess the effect of dataset extension, we next train on the full Slakh2100 dataset ($\textbf{S}_{A}\,$++\,+$\textbf{S}_{B}\,$), resulting in model $\mathcal{T}_{2}\,$. This broader training improves performance on $\textbf{S}_{\text{Full}}\,$and enhances generalization compared to $\mathcal{T}_{1}\,$. However, $\mathcal{T}_{2}\,$still lags behind MSG-LD on other subsets, and its performance on $\textbf{S}_{A}\,$degrades, likely due to increased variability in the training distribution.
@@
The $\mathcal{T}_{3}\,$model, trained on MUSDB18 ($\textbf{M}_{u}\,$) and MoisesDB ($\textbf{M}_{o}\,$) – both containing real recordings rather than synthesized audio – achieves strong performance on their respective test sets. Despite the domain shift, it also performs competitively on Slakh2100, with results comparable to MSDM, highlighting the model’s robustness to cross-domain generalization. Finally, our model trained on the combined dataset – Slakh2100, MUSDB18, and MoisesDB – referred to as $\mathcal{T}_{4}\,$, achieves the best overall results on both $\textbf{M}_{u}\,$and $\textbf{M}_{o}\,$. It also outperforms $\mathcal{T}_{3}\,$on $\textbf{S}_{\text{Full}}\,$, benefiting from the broader training distribution.
@@
We also report results in parentheses, which correspond to mixture generation conditioned on the text prompt “The sound of the bass, drums, guitar, and piano.” For models trained on datasets beyond $\textbf{S}_{A}\,$– such as $\mathcal{T}_{2}\,$, $\mathcal{T}_{3}\,$, and $\mathcal{T}_{4}\,$– the mixtures they generate typically include a broader range of instruments. This can result in a distributional mismatch with the $\textbf{S}_{A}\,$test set, which contains only bass, drums, guitar, and piano, thereby increasing the FAD due to reference-target discrepancy. To mitigate this, we apply text conditioning at inference time using the above prompt to constrain the generated mixture to match the reference instrument set. As shown in the results, this conditioning significantly improves FAD for $\mathcal{T}_{2}\,$, $\mathcal{T}_{3}\,$, and $\mathcal{T}_{4}\,$, yielding performance comparable to the MSDM baseline.
@@
Table 2 presents results for partial generation, evaluated using the sub-FAD metric. This metric computes FAD between the original mixture and a reconstruction formed by combining the given submixture with the generated stems. This evaluation protocol was introduced in prior work on music stem completion [62, 33, 35].
@@
On $\textbf{S}_{A}\,$, our model $\mathcal{T}_{1}\,$performs worse than MSG-LD for single-source imputation, but shows competitive or better performance as the number of generated stems increases. Among our models, $\mathcal{T}_{1}\,$performs best on $\textbf{S}_{A}\,$due to domain alignment. However, for imputation tasks involving broader instrument classes in $\textbf{S}_{B}\,$, models trained on more diverse datasets perform better. Notably, the $\mathcal{T}_{3}\,$model – trained solely on $\textbf{M}_{u}\,$and $\textbf{M}_{o}\,$– achieves strong results on $\textbf{S}_{B}\,$despite not being exposed to Slakh2100 during training. In particular, it performs well on instruments such as organ, pipe, and strings. The fully trained model $\mathcal{T}_{4}\,$further improves performance across several $\textbf{S}_{B}\,$stems, including synth lead, synth pad, and pipe.
@@
Table 2: Partial generation results. Scores are reported using sub-FAD $\downarrow$, which measures the distance between the reference mixture and the sum of given and generated sources. Each column header (e.g., B, D, G) indicates the target source being generated, conditioned on the remaining stems.
@@
MSDM $\times$ $\times$ $\times$ 0.56 1.06 0.49 0.7 2.23 1.56 1.95 1.64 1.83 2.31 3.09 3.53 5.72 3.86 - - - - - - - -
@@
[…]
$\mathcal{T}_{2}\,$$\times$ $\times$ 2.11 2.99 1.99 2.74 4.07 2.32 4.18 3.54 3.9 3.18 4.93 5.69 4.25 4.66 5.96 0.41 1.03 3.66 3.52 2.79 0.88 2.32
@@
Table 3: Source extraction results. Metrics are reported as Log-Mel L1 distance $\downarrow$. For baseline models, scores are shown only for stems included in their fixed output set.
@@
Demucs $\times$ $\times$ $\times$ 1.26 0.92 - - - - - - - - - - 1.59 1.67 1.29 0.89 1.48 1.08 - - - -

### v2

@@ Abstract
[…]
22footnotetext: https://yoongi43.github.io/MGELDM_Samples/
@@ 1 Introduction
Recent advances in generative modeling have significantly accelerated progress in music audio synthesis, inspired by breakthroughs in the language and vision domains. Early autoregressive models such as WaveNet [1] demonstrated the feasibility of end-to-end waveform generation. Since then, two dominant approaches have emerged: (1) discrete-token models, which compress raw audio into quantized codes [2, 3, 4, 5] for sequence modeling [6, 7, 8]; and (2) diffusion-based models [9, 10, 11], which synthesize audio by reversing a noise corruption process in the waveform domain [12, 13]. Building on this foundation, latent diffusion models (LDMs) [14], which operate in a compressed latent space, have delivered substantial gains in synthesis quality and have been successfully applied to music generation tasks [15, 16, 17, 18].
Despite this progress, most music generation models produce a single, mixed waveform, lacking access to the individual instrument stems required for remixing, adaptive arrangement, or downstream production tasks. To recover these components, audio source separation techniques aim to decompose a mixture into its constituent tracks. Discriminative approaches [19, 20, 21, 22] learn to directly regress each source from the input mixture, achieving strong performance. In contrast, generative separation models sample individual sources from a learned prior [23, 24, 25, 26, 27]. Recently, diffusion-based methods have demonstrated strong performance not only in speech separation [28, 29] but also in speech enhancement tasks [30, 31, 32, 33], highlighting their potential for flexible, high-quality source recovery across audio domains.
@@
To address these limitations, we introduce MGE-LDM, a class-agnostic latent diffusion framework that jointly unifies music generation, partial generation, and arbitrary source extraction. Our approach models three interrelated latent variables: mixture, submixture, and source within a single diffusion backbone.
Given a waveform mixture $x^{(m)}$ and a chosen source $x^{(s)}$, we define the submixture as $x^{(u)}=x^{(m)}-x^{(s)}$. A pretrained encoder $E$ maps each waveform to its latent representation:
@@
At inference time, diffusion-based inpainting, conditioned on a subset of $\{z^{(m)},z^{(u)},z^{(s)}\}$, enables class-agnostic partial generation and text-driven source extraction without relying on fixed instrument vocabularies or linear mixing assumptions. A detailed formulation appears in Section 3.
@@
In contrast, our three-track formulation (mixture, submixture, individual source) treats aggregated labels as submixtures. This design integrates them into joint training together with fully separated stems. This dataset flexibility simplifies data collection and enhances model robustness across heterogeneous source structures.
@@ 2 Related Work
Audio Source Separation and Extraction. Audio source separation aims to decompose a polyphonic mixture into its constituent tracks, while source extraction focuses on isolating a particular target sound, often guided by metadata, text prompts, or reference examples. Two dominant paradigms have emerged: discriminative models learn a direct mapping from the input mixture to each target stem via regression losses, either in the waveform domain or in spectrogram representations [41, 19, 42, 20, 21, 22]. In contrast, generative approaches learn probabilistic priors over source distributions and recover individual stems via sampling [23, 24, 26, 25, 43].
@@
Parallel efforts introduced discrete-token models, where audio is encoded into compact code sequences using vector quantization (e.g., VQ-VAE [2]). Jukebox [6] models long-range dependencies over codes using Transformers [60], while recent systems enhance fidelity through residual quantization [3, 4, 5] and hierarchical token modeling, where coarse-to-fine code representations are generated across multiple levels [61, 62, 7, 63]. MusicGen [8] improves decoding efficiency with delayed-token generation, and Instruct-MusicGen [64] extends it for targeted editing via instruction-tuned prompts. In parallel, token-based masked generative modeling techniques, originally developed for vision [65], have been adapted for audio, enabling efficient non-autoregressive synthesis and precise spectrogram inpainting [66, 67].
@@
Multi-Track Music Audio Modeling. Recent studies model multi-track music as a structured composition of interdependent stems. StemGen [75] employs an iterative, non-autoregressive transformer over discrete tokens to generate stems conditioned on text prompts. Jen-1 Composer [76] applies latent diffusion to jointly model four canonical stems (bass, drums, instrument, melody), producing coherent multi-track compositions. MusicGen-Stem [77] combines per-stem vector quantization with an autoregressive decoder to synthesize bass, drums, and aggregated other components, and supports mixture-conditioned accompaniment generation. Diff-A-Riff [78] leverages latent diffusion to co-create stems complementary to given mixtures, later extended with diffusion transformers [79]. Other studies have similarly explored mixture-conditioned stem generation [80, 81].
A separate line of work focuses on joint modeling of synthesis and decomposition within a single diffusion backbone. Multi-Source Diffusion Models (MSDM) [34] model a fixed set of stems (bass, drums, guitar, and piano) within a shared diffusion framework, relying on an additive mixture assumption and a Dirac delta-based posterior sampler, following the EDM formulation for ODE-based sampling [82]. This line of work has since been extended in GMSDI [35], MSG-LD [36], and others [40, 83]. GMSDI enables variable-stem modeling and text-based conditioning but remains grounded in waveform-space additive mixing. MSG-LD adapts latent diffusion for four-stem modeling and classifier-free guidance [84], though it still assumes fixed instrument classes.
@@ 3 Method
Figure 1 presents an overview of the proposed MGE-LDM framework. We train a three-track joint latent diffusion model that learns the distribution over the joint space of mixture, submixture, and source representations, as defined in Eq. (1). During inference, the model performs various tasks by exploiting the inpainting capability of diffusion models. We first describe how training triplets are constructed, then detail how they are used for joint diffusion-based training and inpainting-based inference.
@@ 3.2 Latent Diffusion Training with Three-Track Embeddings
where $z^{(k)}_{0}\in\mathbb{R}^{C\times L}$ are (clean) track embeddings, with $k\in K=\{m,u,s\}$ denoting the track types – mixture, submixture, and source, respectively.
We aim to estimate the score $\nabla_{\textbf{z}_{\tau}}\log q_{\tau}(\textbf{z}_{\tau})$ across continuous noise levels $\tau\in\left[\tau_{\text{min}},1\right]$. To this end, we perturb the clean latent variable $\textbf{z}_{0}$ with Gaussian noise, following:
@@
where the noise scaling coefficients are parameterized as:
@@
[…]
@@ 3.3 Inference via Conditional Sampling in Latent Space
Total Generation. Let $p_{\theta}(z^{(m)},z^{(u)},z^{(s)})$ denote the implicit model distribution whose score we approximate with $f_{\theta}$. To synthesize a complete mixture, we condition only on the mixture prompt embedding $c^{(m)}$ or omit all conditions for unconditional generation. We sample the mixture latent $\hat{z}^{(m)}$ as:
@@
Let $\bar{\mathcal{I}}=\{c_{1},...,c_{J}\}$ be an (ordered) set of CLAP-derived text embeddings, each corresponding to a target source description to be imputed. Let $x^{(u)}_{0}$ denote the waveform mixture of the observed sources, and let $z^{(u)}_{0}=E(x^{(u)}_{0})$ be its latent representation. We initialize the submixture latent with $z^{(u)}_{0}$ and generate each missing source sequentially.
@@
After $J$ iterations, we obtain the full set of imputed sources $\{\hat{z}^{(s)}_{j}\}_{j=1}^{J}$ and reconstruct the final mixture waveform as:
@@ 3.4 Track-aware Inpainting Model with Adaptive Timesteps
where $\alpha_{\tau}$ and $\beta_{\tau}$ is cosine schedule as defined in Eq. (1), following [87]. For notational simplicity, we omit the conditioning vectors $\mathbf{c}$ in this expression. A detailed derivation is provided in Appendix A.1.
Recently, region-aware adaptations of diffusion inpainting, including spatially varying noise schedules [91] and per-pixel timestep conditioning in TD-Paint [92], have demonstrated substantial improvements in semantic consistency by preserving fidelity in observed regions. Inspired by TD-Paint, we extend this idea to three-track music audio by assigning distinct timestep conditions to each track, thereby improving inpainting quality in the latent space.
We describe our track-wise adaptive timestep conditioned model using general notation. Let $K$ be a set of tracks, and let $N=|K|$ be the number of tracks. Define the clean latent tensor and the corresponding noise levels as:
@@
and extend this notation to the cosine noise schedule terms as follows:
@@
The denoiser $f_{\theta}(\textbf{z}_{\boldsymbol{\tau}},\boldsymbol{\tau})$ is trained to regress the velocity target under the v-objective:
@@
[…]
@@
During training, we first sample a noise level $\boldsymbol{\tau}\sim\mathcal{U}([\tau_{\text{min}},1])$ and set the per-track timestep vector $\boldsymbol{\tau}\in\mathbb{R}^{3}$ according to one of the following four patterns:
@@
Under this conditioning strategy, the full-noise setting $\boldsymbol{\tau}=(\tau,\tau,\tau)$ corresponds to learning the standard joint score, as described in Eq. (9). In contrast, a "single-zero" pattern allows the model to learn conditional score functions for the unobserved tracks while treating the others as fixed observations.
For example, when $(\tau_{m},\tau_{u},\tau_{s})=(0,\tau,\tau)$, the model is trained to approximate the gradient:
@@
which the joint-score formulation in Eq. (9) cannot compute in closed form. At inference time, we clamp the observed tracks by setting their noise levels to zero and apply standard reverse diffusion updates to the remaining (missing) tracks. Pseudocode and a detailed theoretical comparison with conventional inpainting methods are provided in Appendix B.
@@ 4 Experimental Setup
Baselines. We use two recent multi-track diffusion models — MSDM [34] and MSG-LD [36] — as baselines, both of which operate on a fixed set of stems: bass, drums, guitar, and piano. In addition to generative and inpainting performance, we assess source extraction capabilities against Hybrid Demucs (HDemucs) [20], which separates the mixture into bass, drums, other, and vocals stems, and AudioSep [53], which performs text-conditioned separation based on natural language queries.
@@
Implementation Details. Our models use the Stable Audio backbone [72], which comprises an autoencoder and a DiT-based diffusion model. To better accommodate per-track variability in the joint latent space, we replace LayerNorm [93] with GroupNorm [94], using three groups to reflect the number of tracks.
@@
All of our models, except the one trained on the full dataset combination ($\textbf{S}_{A}$, $\textbf{S}_{B}$, $\textbf{M}_{u}$, $\textbf{M}_{o}$), are trained for 200K iterations with a batch size of 64, using 16 kHz audio segments of 10.24 seconds. The full combination model is trained for 320K iterations with a batch size of 128. During sampling and inpainting, we apply classifier-free guidance (CFG) with a guidance scale of 2.0 and a per-track dropout probability of $p=0.1$. All diffusion-based samples – including those from baseline models – are generated using 250 inference steps. We adopt DDIM sampling [96] for all our models, while each baseline uses its originally proposed sampling method.
5 Results
Table 1: Total generation results. Reported scores are FAD $\downarrow$, computed against mixture references from each test set. Values in parentheses indicate generation results conditioned on the prompt "The sound of the bass, drums, guitar, and piano", as detailed in Section 5.1. Among our models, $\mathcal{T}_{1}$ provides a fair comparison with baseline methods, as it is trained on the same dataset, while $\mathcal{T}_{2}$$\mathcal{T}_{4}$ illustrate the effects of progressive dataset scaling. Bold values indicate the best results in each column, and underlined values denote the best results among models trained on the same $\textbf{S}_{A}$ set.
@@ 5.1 Music Generation
To assess the effect of dataset extension, we next train on the full Slakh2100 dataset ($\textbf{S}_{A}$$\textbf{S}_{B}$), resulting in model $\mathcal{T}_{2}$. This broader training improves performance on $\textbf{S}_{\text{Full}}$ and enhances generalization compared to $\mathcal{T}_{1}$. However, $\mathcal{T}_{2}$ still lags behind MSG-LD on other subsets, and its performance on $\textbf{S}_{A}$ degrades, likely due to increased variability in the training distribution.
@@
The $\mathcal{T}_{3}$ model, trained on MUSDB18 ($\textbf{M}_{u}$) and MoisesDB ($\textbf{M}_{o}$), both containing real recordings rather than synthesized audio, achieves strong performance on their respective test sets. Despite the domain shift, it also performs competitively on Slakh2100, with results comparable to MSDM, highlighting the model’s robustness to cross-domain generalization. Finally, our model trained on the combined dataset comprising lakh2100, MUSDB18, and MoisesDB (denoted as $\mathcal{T}_{4}$), achieves the best overall results on both $\textbf{M}_{u}$ and $\textbf{M}_{o}$. It also outperforms $\mathcal{T}_{3}$ on $\textbf{S}_{\text{Full}}$, benefiting from the broader training distribution.
@@
We also report results in parentheses, which correspond to mixture generation conditioned on the text prompt “The sound of the bass, drums, guitar, and piano.” For models trained on datasets beyond $\textbf{S}_{A}$, including $\mathcal{T}_{2}$, $\mathcal{T}_{3}$, and $\mathcal{T}_{4}$, the generated mixtures generally contain a wider variety of instruments. This can result in a distributional mismatch with the $\textbf{S}_{A}$test set, which contains only bass, drums, guitar, and piano, thereby increasing the FAD due to reference-target discrepancy. To mitigate this, we apply text conditioning at inference time using the above prompt to constrain the generated mixture to match the reference instrument set. As shown in the results, this conditioning significantly improves FAD for $\mathcal{T}_{2}$, $\mathcal{T}_{3}$, and $\mathcal{T}_{4}$, yielding performance comparable to the MSDM baseline.
@@
Table 2 presents results for partial generation, evaluated using the sub-FAD metric. This metric computes the FAD between the original mixture and a reconstruction formed by combining the given submixture with the generated stems. This evaluation protocol was introduced in prior work on music stem completion [63, 34, 36].
@@
On $\textbf{S}_{A}$, our model $\mathcal{T}_{1}$performs worse than MSG-LD for single-source imputation but shows competitive or better performance as the number of generated stems increases. Among our models, $\mathcal{T}_{1}$ performs best on $\textbf{S}_{A}$ due to domain alignment. However, for imputation tasks involving broader instrument classes in $\textbf{S}_{B}$, models trained on more diverse datasets perform better. Notably, the $\mathcal{T}_{3}$ model, trained solely on $\textbf{M}_{u}$ and $\textbf{M}_{o}$, achieves strong results on $\textbf{S}_{B}$ despite not being exposed to Slakh2100 during training. In particular, it performs well on instruments such as organ, pipe, and strings. The fully trained model $\mathcal{T}_{4}$ further improves performance across several $\textbf{S}_{B}$ stems, including synth lead, synth pad, and pipe.
@@
Table 2: Partial generation results. Scores are reported using sub-FAD $\downarrow$, which measures the distance between the reference mixture and the sum of given and generated sources. Each column header (e.g., B, D, G) indicates the target source being generated, conditioned on the remaining stems. Bold and underlined values follow the same convention as in Table 1.
@@
MSDM $\times$ $\times$ $\times$ 0.56 1.06 0.49 0.70 2.23 1.56 1.95 1.64 1.83 2.31 3.09 3.53 5.72 3.86 - - - - - - - -
@@
MGE (ours) $\mathcal{T}_{1}$ $\times$ $\times$ $\times$ 1.02 1.41 1.17 1.19 1.15 1.29 1.25 1.69 1.65 1.14 1.80 1.84 1.45 1.84 1.45 0.68 0.23 3.48 5.58 1.38 4.47 1.08
$\mathcal{T}_{2}$ $\times$ $\times$ 2.11 2.99 1.99 2.74 4.07 2.32 4.18 3.54 3.90 3.18 4.93 5.69 4.25 4.66 5.96 0.41 1.03 3.66 3.52 2.79 0.88 2.32
@@
Table 3: Source extraction results. Metrics are reported as Log-Mel L1 distance $\downarrow$. For baseline models, scores are shown only for stems included in their fixed output set. Bold and underlined values follow the same convention as in Table 1.
@@
HDemucs $\times$ $\times$ $\times$ 1.49 0.90 - - - - - - - - - - 1.50 1.99 1.53 0.83 1.71 1.10 - - - -